• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Moffit

    Log entry time 00:31:20 on September19,2005

    Entry number 155148

    keyword=Hall-C using Hall-B laser. Feedback on hall-C off

    Hall-C laser lost it's lock.  MCC is now using the Hall-B diode laser until the 
    problem can be solved tomorrow morning by the source group.
    
    The hall-C IA is useless until the fix, so I've turned off CIA feedback and set 
    it to 5.05 (off).
    
    An expert (Lisa, Ryan, or Bryan) will need to turn this back on (in the 
    panFFB.db file).
    
    Hall-C charge asymmetry is currently -1500ppm.